# Contact SLAM: An Active Tactile Exploration Policy Based on Physical Reasoning Utilized in Robotic Fine Blind Manipulation Tasks

### v1

@@ I Introduction
In robotic manipulation tasks, vision plays an irreplaceable role. Through visual servoing, a robot can identify the target of manipulation and control its manipulator to complete the task [1]. In this process, the tactile or force sensing is mainly used to interact with the environment after the vision has already provided the target position, leading the manipulator to achieve its operational goal with the dual guidance of vision and force sensing [2].
However, in certain tasks, when vision fails or is occluded, the robot can no longer obtain real-time scene state information via visual feedback. We refer to such manipulation tasks as “blind manipulation” tasks, in which the robot must rely solely on tactile and force information. Consequently, researchers have begun investigating how to transform contact data into state information about the scene [3, 4]. However, unlike directly contacting and measuring the environment, in general-purpose manipulation scenarios, the end-effector is often equipped with a two-finger gripper or a dexterous hand to handle different objects, and then uses these objects to interact with the environment, which means the force–tactile information captured cannot be directly used to determine the environment’s state. Although some studies have investigated the perception process during manipulation with grippers holding objects [5, 6], they typically stop at the perception stage without conducting an in-depth study on task completion.
During human manipulation, in addition to visual awareness of the process, there is also contact awareness. By judging the force sensed at the fingertips, people can estimate whether the manipulation goal has been reached and, if not, how far they are from the target state. In robotic manipulation, visual servoing generally determines the target object in the scene through active perception and cognition; however, in blind manipulation tasks, this mechanism is absent. Some researches have tried to transform the force or tactile signals into some forms of contact state [7, 8], but these methods usually have strong data dependency. Inspired by human manipulation, we introduce physically-driven contact cognition into robotic blind manipulation. By reasoning about the contact state between the grasped object and the environment from contact signals, the robot can continuously estimate and explore the environment until ultimately completing the manipulation task, as illustrated in Fig. 1.
@@
    A tactile perception method for estimating the force and its position acting on the grasped object is proposed. Given the precise geometric shape of the grasped object and the motion trends of tactile sensors mounted on the two-finger gripper, the proposed method can determine the forces exerted on the grasped object, as well as the contact position.
@@
    A contact-based SLAM method for environment perception and localization is proposed. This method further develops and extends the scope of indirect perception and tactile SLAM, enabling the perception and understanding of the scene state based on prior scene knowledge and real-time tactile feedback.
@@
    A planning method for fine blind manipulation based on active tactile exploration is proposed. This approach balances environmental exploration with task completion efficiency, plans optimal action strategies according to the current understanding of the environment, and reduces the uncertainties of the manipulation scene gradually.
II Related Work
II-A Tactile Perception During Manipulation Process
Compared with visual sensing, tactile sensing provides more comprehensive information about the contact state with the environment during manipulation, including the geometry of the contacted object[9], the contact force/torque and its distribution[10], and contact event triggering[11]. With these perceptual features, they can be further applied to robotic manipulation tasks[12, 13]. In terms of interaction with the environment, end-to-end policy generation is typically implemented through a sim-to-real approach[14]. Alternatively, tactile sensing can be used as an independent observer[15], whose results are then fed into the control loop.
II-B Localization and Mapping Method based on Tactile Sensing
Sudharshan Suresh et al.[16] proposed using force–tactile sensing to simultaneously track an object’s position and model its contour, a method they termed Tactile SLAM. Jialiang Zhao et al.[17] introduced the concept of FingerSLAM. Paloma Sodhi et al.[18] applied the SLAM modeling concept to object manipulation control, achieving tactile-based pushing operations. Daolin Ma et al.[5] proposed a method that uses visuo-tactile sensor to estimate the relative pose of an object grasped by a two-finger gripper, as well as the location of external contact points and contact lines. Building on this idea, Sangwoon Kim et al.[6] proposed a method for estimating the state of an object during manipulation to maintain its stability, along with an approach for perceiving external contact lines and enabling general peg-in-hole assembly[19].
II-C Cognition and Reasoning During Manipulation Process
Object cognition primarily focuses on visual sensing. Chaofan Zhang et al. proposed VTLA [20], which integrates tactile modality into a large-scale manipulation model. Features representing contact constraints are generally obtained by two approaches: one is through force–tactile sensors, and the other is through a world model [21]. State cognition tends to focus on low-level state feedback and control, such as determining whether specific signals relevant to task completion are present during contact [22] or designing contact feedback control loops [23, 24] to replicate particular contact signals.
III Task Formulation
In blind manipulation tasks, several kinds of objects need to be noticed. The grasped objects, which are grasped by the gripper, and the tactile sensors could only perceive the tactile signals between the grasped objects and the gripper. The manipulated objects, which are used to interact with the grasped objects, including unmovable obstacles and movable objects. The last is the target region, which the grasped objects or manipulated objects need to reach. The detailed scenes are shown in section V-A.
@@ III Task Formulation
IV Methods
IV-A Contact Force and Pose Perception of the Grasped Object
[…]
When calculating the torque, the reference point must be chosen first. The measured resultant torques are calculated in the left and right sensor coordinate frames; however, what we need is the resultant torque calculated in the gripper coordinate frame about the equivalent rotation point: $O_{e}$. By analyzing the torque caused by a single force on a single point, the relationship between force $(f^{x}_{i},f^{y}_{i},f^{z}_{i})$ and torque $(m^{x}_{i},m^{y}_{i},m^{z}_{i})$ is as follows:
$\begin{bmatrix}m^{x}_{i}\\ m^{y}_{i}\\ m^{z}_{i}\end{bmatrix}=\begin{bmatrix}0&-f^{z}_{i}&f^{y}_{i}\\ f^{z}_{i}&0&-f^{x}_{i}\\ f^{y}_{i}&-f^{x}_{i}&0\end{bmatrix}\cdot\begin{bmatrix}O_{x}\\ O_{y}\\ O_{z}\end{bmatrix}$ (2)
where $(O_{x},O_{y},O_{z})$ is the pose of the force application point. Thus, when the reference point changes from $O_{1}$ to $O_{2}$ and doesn’t consider the coordinate system rotation transformation, the torque changes to $((m^{x}_{i})_{O_{2}},(m^{y}_{i})_{O_{2}},(m^{z}_{i})_{O_{2}})$:
$\begin{bmatrix}(m^{x}_{i})_{O_{2}}\\ (m^{y}_{i})_{O_{2}}\\ (m^{z}_{i})_{O_{2}}\end{bmatrix}=\begin{bmatrix}0&-f^{z}_{i}&f^{y}_{i}\\ f^{z}_{i}&0&-f^{x}_{i}\\ f^{y}_{i}&-f^{x}_{i}&0\end{bmatrix}\cdot\begin{bmatrix}O^{x}_{1}+\Delta X^{O_{2}}_{O_{1}}\\ O^{y}_{1}+\Delta Y^{O_{2}}_{O_{1}}\\ O^{z}_{1}+\Delta Z^{O_{2}}_{O_{1}}\end{bmatrix}$ (3)
where $\Delta X^{O_{2}}_{O_{1}},\Delta Y^{O_{2}}_{O_{1}},\Delta Z^{O_{2}}_{O_{1}}$ are vectors from point $O_{2}$ to $O_{1}$.
In this way, the resultant torques applied to the grasped object about the equivalent rotation point $O_{e}$ could be written as follows:
$\begin{split}\begin{bmatrix}(M_{x})_{O_{e}}^{L/R}\\ (M_{y})_{O_{e}}^{L/R}\\ (M_{z})_{O_{e}}^{L/R}\end{bmatrix}=\begin{bmatrix}(M_{x})^{L/R}\\ (M_{y})^{L/R}\\ (M_{z})^{L/R}\end{bmatrix}+&\begin{bmatrix}0&-F_{z}^{L/R}&F_{y}^{L/R}\\ F_{z}^{L/R}&0&-F_{x}^{L/R}\\ F_{y}^{L/R}&-F_{x}^{L/R}&0\end{bmatrix}\\ &\cdot\begin{bmatrix}\Delta X^{O_{e}}_{O_{L/R}}\\ \Delta Y^{O_{e}}_{O_{L/R}}\\ \Delta Z^{O_{e}}_{O_{L/R}}\end{bmatrix}\end{split}$ (4)
where $F_{i}^{L},M_{i}^{L}$ and $F_{i}^{R},M_{i}^{R}$ are corresponding force and torque components ($i=x,y,z$) in the frame of the left sensor and right sensor. The parameters need to be measured are $\Delta X^{O_{e}}_{O_{L/R}},\Delta Y^{O_{e}}_{O_{L/R}},\Delta Z^{O_{e}}_{O_{L/R}}$.
Using the computed force and torque values in Eq. 1 and 4, we can apply the above equations to determine the force application region $(C_{x},C_{y},C_{z})$ of the grasped object:
$\begin{bmatrix}(M_{x})_{O_{e}}^{L}+(M_{x})_{O_{e}}^{R}\\ (M_{y})_{O_{e}}^{L}+(M_{y})_{O_{e}}^{R}\\ (M_{z})_{O_{e}}^{L}+(M_{z})_{O_{e}}^{R}\end{bmatrix}=\begin{bmatrix}0&-F_{z}&F_{y}\\ F_{z}&0&-F_{x}\\ F_{y}&-F_{x}&0\end{bmatrix}\cdot\begin{bmatrix}C_{x}\\ C_{y}\\ C_{z}\end{bmatrix}$ (5)
IV-B Contact SLAM Analysis
@@
    where $B(\cdot)$ means the boundary of the objects or obstacles’ position. The parameter $P_{t}$ is the distribution range of particles at the current time, which is shown in section IV-C. Initially, the $e_{0}$ is set to a range that could cover the potential position of interested objects.
@@
IV-C Active Tactile Exploration Policy
In the process of blind manipulation, the exact pose of the object is unknown. To achieve the goal of estimating relative contact regions, the system must generate exploratory actions to interact with the environment, continuously estimate the contact state during the contact process, and adjust the action based on the tactile sensor feedback. Based on these, we draw inspiration from active localization methods[25] and propose the Active Tactile Exploration Policy(ATEP) method. The detailed workflow is shown as follows:
@@ IV-C Active Tactile Exploration Policy
Step 2, Information gain evaluation: For each potential motion direction of the object, the expected information gain is calculated, which is represented as the types of contact states and the distance variance of motion distance. The direction with the maximum information gain is then selected as the movement direction. The detailed process is in Algorithm 1.
@@
5:for $a$ in $range(A)$ do
6:  for $p$ in $range(Pose)$ do
7:   Calculate the contact state: z_predz\_pred at $p$ with action $a$;
8:   Add z_predz\_pred to $Z_{a}$;
9:   Calculate the distance $d$ to get the state z_predz\_pred;
10:   Add $d$ to $D_{a}$;
11:  end for
12:end for
13: Choose the action which could get the most information: $\max_{a}\{\alpha_{1}\cdot entropy(Z_{a})+\alpha_{2}\cdot variance(D_{a})\}$.
@@
Step 4, Contact pair selection and updation of particles: After detecting a tactile signal change, according to the method described in SectionIV-A, we can estimate the direction of the force acting on the grasped object. From the perspective of object dynamics, the force direction and the outward normal vector of the object’s contour satisfy the following relationship: $F_{G}^{t}\cap n_{i}>0,F_{G}^{t}\cap n_{j}<0$. Thus, the contact pairs can be represented as: $(i,j)=\{n_{i}=-n_{j}\}$, and the distribution of particels is updated as follows: $P_{t}=P_{t-1}\cap(B(S_{e_{i}})\cup B(S_{l_{j}}))$.
@@
3:if $len(P_{t})<30$ then
4:  Add more particles to the scene;
5:end if
6: Calculate the local optimal particles: $peaks=\{p_{i}^{t}|p_{i}^{t}\in P_{t},w_{i}^{t}>1/len(P_{t}),i=1,2,...n\}$;
7:if $len(peaks)>1$ then
8:  Information gain evaluation and obtain action $\pi$;
9:  Execute $\pi$ and monitor force until the contact occurs;
10:  Contact pair selection: $P_{t+1}=P_{t}\cap(B(S_{e_{i}})\cup B(S_{l_{j}}))$;
11:  Update the particles’ weights;
12:  return to “Calculate the local optimal particles”.
13:end if
14:if $len(peaks)<n_{thr}$ and $B(peaks)<\delta_{thr}$ then
15:  Find the optimal trajectory $\pi$;
16:  Execute $\pi$, until the alignment factor satisfies $\theta_{ali}=1$.
17:end if
@@
V Experimental Studies
V-A Experimental Setup
@@
The first task scenario involves a blind peg-in-hole assembly of a power socket component of two different standards. The robot is required to grasp the socket and insert it into the correct receptacle. The second task scenario is a blind pushing task in the presence of obstacles. In this scenario, the operator has prior knowledge of the approximate position of the block to be pushed, the target location, as well as the geometrical shapes of the obstacles present in the environment. However, the exact positions of these obstacles within the scene are unknown.
V-B Contact Point Localization of the Grasped Object
For tactile perception, we first perform an unknown-parameter estimation process, and then use the estimated parameters together with Eq. 5 to verify the effectiveness of the proposed contact position prediction method.
In the parameter estimation stage, we design the experimental setup shown in Fig. 4(a). In this setup, a two-finger gripper equipped with Tac3D sensors clamps a calibration block, and each face is uniformly divided into several grooves along both horizontal and vertical directions. During the experiment, several points are selected on each face, and a contact force is applied at each selected point. Using the method described in Eq. 1, we compute the three-dimensional contact force exerted at each point and extract the resultant moment applied at the origin of each sensor coordinate system.
During the data collection stage, the contact point coordinates $(C_{x},C_{y},C_{z})$ in the object coordinate system are assumed to be known. We then adopt the least-squares method to solve for the unknown parameters $\Delta X^{O_{e}}_{O_{L/R}},\Delta Y^{O_{e}}_{O_{L/R}},\Delta Z^{O_{e}}_{O_{L/R}}$. In the data validation stage, given new input force and moment data, the estimated parameters are substituted into Eq. 4 and 5 to calculate the coordinates of the contact point. The prediction results of the contact points in the training set are shown in Fig. 5.
Figure 5: The predicted contact points vs the actual contact points.
As illustrated in Fig. 5, the fitting errors remain at a very small magnitude, with most errors concentrated around zero and the maximum error not exceeding 0.5mm. These experimental results demonstrate that the proposed tactile-based contact position estimation method is both accurate and effective.
V-C Results of the Socket Assembly Experiments

### v2

@@ I Introduction
In robotic manipulation tasks, vision plays an irreplaceable role. Through visual servoing, a robot can identify the target of manipulation and control its manipulator to complete the task [9]. In this process, the tactile or force sensing is mainly used to interact with the environment after the vision has already provided the target position [20].
However, in certain tasks, when vision fails or is occluded, the robot can no longer obtain real-time scene state information via visual feedback. We refer to such manipulation tasks as “blind manipulation” tasks, in which the robot must rely solely on tactile and force information. Compared with visual sensing, tactile sensing provides more comprehensive information about the contact state with the environment during manipulation, including the geometry of the contacted object[17], the contact force/torque and its distribution[21], and contact event triggering[19]. Some researches have tried to transform the force or tactile signals into some forms of contact state [10, 7, 5], The manipulation based on contact states tends to focus on low-level state feedback and control, such as determining whether specific signals relevant to task completion are present during contact[12] or designing contact feedback control loops[4, 18, 6] to replicate particular contact signals, but these methods usually have strong data dependency, and the utilized scene is constrained strictly. With these perceptual features, they can be further applied to robotic manipulation tasks[3, 23]. In terms of interaction with the environment, end-to-end policy generation is typically implemented through a sim-to-real approach[2]. Although some studies have investigated the perception process during manipulation with tactile sensing, they typically stop at the perception stage without conducting an in-depth study on task completion. The deatiled comparison between the visual and tactile servo control process is shown in Fig. 1
@@
During human manipulation, in addition to visual awareness of the process, there is also contact awareness. By judging the force sensed at the fingertips, people can estimate whether the manipulation goal has been reached and, if not, how far they are from the target state. This is a typical simultaneously localization and mapping(SLAM) problem. Sudharshan Suresh et al.[16] proposed using force–tactile sensing to simultaneously track an object’s position and model its contour, a method they termed Tactile SLAM. Jialiang Zhao et al.[22] introduced the concept of FingerSLAM. Paloma Sodhi et al.[15] applied the SLAM modeling concept to object manipulation control, achieving tactile-based pushing operations. However, in blind manipulation tasks, except for the missed viisual information, there also exists uncertainties, which could prevent the task to be finished, but the existing tactile methods don’t consider this.
In manipulation planning process, the uncertainty eliminating is widely used in the contact process[14, 13], but they usually use the binary force singal, and they often cosider one target object without interference from other objects. So Inspired by tactile SLAM methods and contact-driven blind manipulation methods, we introduce physically-driven contact cognition into robotic blind manipulation. By reasoning about the contact state between the grasped object and the environment from contact signals, the robot can continuously estimate and explore the environment until ultimately completing the manipulation task, as illustrated in Fig.1.
@@
    A contact state-based pipeline for executing the contact-rich manipulation is proposed. The contact-rich manipulation process is defined as the contact state translation process, and the relative pose between the grasped object and scene could be inferred from the contact state.
@@
    A contact event-based SLAM method for environment perception and localization is proposed. This method defines the contact event and further develops and extends the scope of tactile SLAM, proposing a post-estimation framework for relative pose estimation, enabling the perception and understanding of the scene state based on prior scene knowledge and contact event feedback.
@@
    An effective planning method for fine blind manipulation based on active tactile exploration is proposed. This approach balances environmental exploration with task completion efficiency, plans optimal action strategies according to the current understanding of the environment, and reduces the uncertainties of the manipulation scene gradually.
II Problem Formulation: Contact SLAM Analysis
In blind manipulation tasks, several kinds of objects need to be noticed. The grasped objects, which are grasped by the gripper, and the tactile sensors could only perceive the tactile signals between the grasped objects and the gripper. The manipulated objects, which are used to interact with the grasped objects, including unmovable obstacles and movable objects. The last is the target region, which the grasped objects or manipulated objects need to reach. The detailed scenes are shown in section IV-A.
@@ II Problem Formulation: Contact SLAM Analysis
III Methods
III-A Contact Force Estimation Analysis
[…]
III-B State Perception and Optimization
@@
    where $B(\cdot)$ means the boundary of the objects or obstacles’ position. The parameter $P_{t}$ is the distribution range of particles at the current time, which is shown in section III-C. Initially, the $e_{0}$ is set to a range that could cover the potential position of interested objects.
@@
III-C Contact Localization and Active Tactile Exploration Policy
In the process of blind manipulation, the exact pose of the object is unknown. To achieve the goal of estimating relative contact regions, the system must generate exploratory actions to interact with the environment, continuously estimate the contact state during the contact process, and adjust the action based on the tactile sensor feedback. Based on these, we draw inspiration from active localization methods[1] and propose the Contact Localization and Active Tactile Exploration Policy method. The detailed workflow is shown as follows:
@@ III-C Contact Localization and Active Tactile Exploration Policy
Step 2, Information gain evaluation: For each potential motion direction of the object, the expected information gain is calculated, which is represented as the types of contact states of motion distance. We choose the entropy as the justance criterion, and calculate the entropy of each action as follows:
$entropy(Z_{a})=\sum_{z\in Z_{a}}-p(z)\log p(z)$ (8)
Where $Z_{a}$ is the predicted contact state set of all particles with the action $a$.
Afer getting the information gain of each potential action, we choose the action which could get the most information as the moving direction: $\max_{a}\{entropy(Z_{a})\}$.
Step 4, Contact pair selection and updation of particles: After detecting a tactile signal change, we can estimate the direction of the force acting on the grasped object. From the perspective of object dynamics, the force direction and the outward normal vector of the object’s contour satisfy the following relationship: $F_{G}^{t}\cap n_{i}>0,F_{G}^{t}\cap n_{j}<0$. Thus, the contact pairs can be represented as: $(i,j)=\{n_{i}=-n_{j}\}$, and the distribution of particels is updated as follows: $P_{t}=P_{t-1}\cap(B(S_{e_{i}})\cup B(S_{l_{j}}))$.
After updating the weights of particles, the local optimum detection process is repeated, and the particles are further updated: $P_{t}=peaks$.
If particles’ distribution is less than the predefined threshold, the process turns to Step 7 to find the optical policy to finish the task. If not, the process turns to Step 6 to resample more particles.
Step 6, Resample particles: After updating the weights of all particles, we check the number of particles. If the number of particles is less than a predefined threshold (e.g., 50), we resample additional particles within the current potential region to ensure sufficient particle density for accurate localization. Based on the particle filter method, we choose to resample the same number of particles each iteration.
Return to the Step 1 and continue the exploration process.
Step 7, The generation of final manipulation policy: When the particles’ distribution is smaller than the set threshold, the manipulation policy is generated. The optimal policy is represented as moving from the estimated object pose to the target pose. In the motion planning process, we choose the A_star algorithm as the planning method.
III-D The improved tips of Active Exploration
In order to improve the efficiency of the active exploration, we adopt the following improvements:
First, the distance variance criterion is added to the process of calculating the information gain. When the distribution of particles is small, the entropy calculation could become very small, too. That means the information gain is useless in the choice of action direction. When the entropy calculation is small, the distance variance is added to the information gain. A larger distance criterion value indicates greater variability in the motion of the current particle distribution, enabling more particles to be distinguished based on the discrepancy between the estimated motion distance and the actual motion distance.
The deatils of Gain information calculation is shown in Algorithm 1.
@@ III-D The improved tips of Active Exploration
5:if $len(P_{t})=1$: then
6:  Choose the optical trajectory;
7:end if
8:if $len(P_{t})>1$: then
9:  for $a\in A$ do
10:   for $p\in P_{t}$ do
11:    Calculate the contact state: z_predz\_pred at $p$ with action $a$;
12:    Add z_predz\_pred to $Z_{a}$;
13:    Calculate the distance $d$ to get the state z_predz\_pred;
14:    Add $d$ to $D_{a}$;
15:   end for
16:   Calculate the entropy of $Z_{a}$: $entropy(Z_{a})=\sum_{z\in Z_{a}}-p(z)\log p(z)$;
17:   Calculate the variance of $D_{a}$: $var(D_{a})=\sum_{d\in D_{a}}(d-\mu(D_{a}))^{2}$;
18:  end for
19:  if $\max\{entropy(Z_{a})\}<e_{thr}$: then
20:   Choose the action based on the Distance variance: $\max_{a}\{var(D_{a})\}$;
21:  else
22:   Choose the action which could get the most information: $\max_{a}\{entropy(Z_{a})\}$.
23:  end if
24:end if
@@
Second, we adjust the number of particles to resample. Instead of resampling the same particles in each episode. In environments with imperfect or imprecise modeling, overly rapid particle depletion may cause some plausible distribution regions to be discarded prematurely. A straightforward solution is to increase the number of particles during resampling. However, if the resampled particle count is fixed, an insufficient number of particles fails to improve localization accuracy, while an excessive number leads to particle overcrowding in later iterations, or even to a continuous increase in the total particle count during the filtering process. To balance exploration efficiency and localization accuracy, we therefore design an adaptive particle resampling mechanism. This mechanism maintains a reasonable particle density in all potentially valid distribution regions, thereby reducing the probability that correct hypotheses are eliminated due to incidental or unexpected events.
@@
3: Local optimum detection: $P_{t}=peaks=\{p_{i}^{t}|p_{i}^{t}\in P_{t},w_{i}^{t}>1/len(P_{t}),i=1,2,...n\}$;
4: Information gain evaluation and obtain action $\pi$;
5: Execute $\pi$ and monitor force until the contact occurs;
6: Updation of particles: $P_{t+1}=P_{t}\cap(B(S_{e_{i}})\cup B(S_{l_{j}}))$;
7: Update the particles’ weights: $w_{i}^{t+1}=w_{i}^{t}\cdot\mathcal{P}(z_{obs}^{t},z_{pred}(p_{i}-(T-t)/T\cdot Distance))$, $W_{t+1}={w_{i}^{t+1}|i=1,2,...,len(P_{t+1})}$;
8: Resample particles: $P_{t+1}^{\prime}=AR(P_{t+1},P_{t},W_{t+1})$
9:if $B(peaks)>\delta_{thr}$ then
10:  return to “Local optimum detection”.
11:else
12:  Find the optimal trajectory $\pi$;
13:  Execute $\pi$, until the alignment factor satisfies $\theta_{ali}=1$;
14:end if
@@
IV Experimental Studies
IV-A Experimental Setup
@@
The first task scenario involves a blind peg-in-hole assembly of a power socket component of two different standards. The robot is required to grasp the socket and insert it into the correct receptacle. The second task scenario is a blind pushing task in the presence of obstacles. In this scenario, the operator has prior knowledge of the approximate position of the block to be pushed, the target location, as well as the geometrical shapes of the obstacles present in the environment. However, the exact positions of these obstacles within the scene are unknown. The chosen two tasks are all typical blind indirect contact manipulation tasks, because the contact with the scene needs to be referred from the contact signals between the gripper and the grasped object.
  • $\bullet$
    Evaluation metrics: After the mapping process, we calculate the localization error represented as root mean square error(RMSE) and iteration times. After the manipulation process, we evaluate the whole exploration time of manipulation, as well as the success rate.
  • $\bullet$
    Compared methods: We choose the Spiral-hole-searching method, which is a typical passive searching method, as the baseline and control method.
IV-B Results of the Socket Assembly Experiments